OpenStudy (anonymous):

Can someone show me how to graph sin(x * pi/2) step by step?

OpenStudy (anonymous):

Rather \[f(x) = (x \times \pi/2)\]

OpenStudy (anonymous):

Rather \[f(x) = \sin [x(\pi/2)]\]

OpenStudy (anonymous):

y=Asin(B(x-C))+D

OpenStudy (anonymous):

Ok so you are defining the coefficients, I get it so far, continue =0)

OpenStudy (anonymous):

I think...

OpenStudy (anonymous):

haha okay! Good so you you know what a sin(x) looks like? I just want to know what you know and dont know.

OpenStudy (anonymous):